# Multi-modal graph neural networks for localized off-grid weather forecasting

### v1

@@ 1 Introduction
  • We evaluate our GNN against a range of data-driven and non-data-driven off-grid weather forecasting methods. Our model achieves significantly better performance. It decreases the MSE by 55.22% comparing to the best performing MLP model which reduces the MSE by 82.55% comparing to interpolated ERA5.
  • We conducted an ablation of ERA5 inputs and observed that a GNN with ERA5 nodes achieves half the MSE of a GNN without ERA5, indicating that—even in the presence of historical station data—global atmospheric dynamics inform local weather patterns.
@@ 2 Related Work
Weather forecasting has long been a challenging problem in atmospheric sciences, with efforts dating back centuries. Since the advent of numerical weather prediction (NWP) in the mid-20th century, most forecast simulations have been conducted on a regular grid, dividing the atmosphere into evenly spaced discrete points to solve complex partial differential equations. This grid-based approach has remained the foundation of many numerical weather forecasting models such as the Integrated Forecast System (ECMWF, 2022) and High-Resolution Rapid Refresh (Dowell et al., 2022). In recent years, machine learning (ML) has gained traction as a promising tool in weather forecasting (Bauer et al., 2015), offering new techniques to improve accuracy and computational efficiency. These ML weather models can be roughly divided into two categories: end-to-end models and foundation models. FourCastNet (Pathak et al., 2022), GraphCast (Lam et al., 2023), Pangu-Weather (Bi et al., 2023), AIFS (Lang et al., 2024) and NeuralGCM (Kochkov et al., 2024) are end-to-end models trained directly to make weather forecasts. In contrast, AtmoRep (Lessig et al., 2023), ClimaX (Nguyen et al., 2023), Aurora (Bodnar et al., 2024) and Prithvi WxC (Schmude et al., 2024) are foundation models that are first trained with a self-supervision task and then fine-tuned for weather forecasting. However, the training data for ML models largely stem from traditional gridded numerical simulations such as ERA5 (Hersbach et al., 2020). As a result, the ML models themselves still typically maintain the grid-based paradigm even within their more modern forecasting approach. One major disadvantage of gridded weather forecasting is that it is usually limited by its fixed resolution such that it cannot accurately reflect fine-grained local weather patterns (although efforts towards limited area modeling have recently been made (Oskarsson et al., 2023)). Other works focusing on increasing the forecast resolution (Harder et al., 2023; Yang et al., 2023; Prasad et al., 2024) exist, but their methods are mostly tested on synthetic datasets. In this work, we propose a multi-modal graph neural network which can effectively downscale gridded weather forecasting to match real-world local weather dynamics.
@@ 4 Results
We report the model performance for the 2023 test set as the mean squared error (MSE) of the predicted wind vector against the actual wind vector (equivalently, mean sum of squared errors in $u$ and in $v$). Figure 6 and Table A.4 summarize the MSE for the different lead times and models. The total MSE for each experiment (summed across lead times) is also reported in Table A.4.
@@
Interpolated ERA5 data does not describe the local wind conditions accurately, with an MSE at all lead times of 9.6 $m^{2}/s^{2}$.
Figure 6: Mean squared error (MSE) of the wind vector for different models calculated over all test set time steps and stations, from 1 hour to 48 hours lead times. The interpolated ERA5 error is calculated by linearly interpolating ERA5 to each weather station; it is the error between the global and local dataset. The persistence is calculated by offsetting the ground truth (MADIS) by the lead time. The results for the multi-layer perceptron (MLP) and message passing neural network (MPNN) are reported with and without the inclusion of ERA5 as input. All methods outperform the ERA5 interpolation, all ML methods outperform the persistence, and the MPNN + ERA5 outperforms all other methods.
(Note that the MSE for interpolated ERA5 is constant across lead times, because the “future” values of ERA5 are assumed to be known. This is the best case for a gridded NWP forecast or an ML weather model trained to predict ERA5; if ERA5 were replaced by the output of a global forecasting model, the error would increase with lead time.) Spatially, when considering the total MSE for each station (Figure 7), ERA5 appears to consistently misrepresent local weather patterns. The largest magnitude errors are generally concentrated along the coast, due to the higher average wind speeds there, but the highest relative errors are inland.
The persistence baseline shows a low MSE for short lead times—lead times 1, 2, and 4 hours all have MSEs below 1.6 $m^{2}/s^{2}$. This is expected due to temporal auto-correlation. MSE rapidly increases with longer lead times, reaching a maximum MSE of 5.6 $m^{2}/s^{2}$ at 36 hours. Hour 48 shows a slight decrease in MSE, hinting at some daily periodicity in wind patterns. Spatially, most of the high MSE is also concentrated along the coast, with some exceptions inland. The MSE of inland stations is on average substantially lower for the persistence than the interpolated ERA5 baseline.
@@
The MLP + ERA5 method significantly improves on the non-ML methods, showing an MSE below 2.6 $m^{2}/s^{2}$ for all lead times. The MLP error increases with lead hours as expected, reaching the beginning of a plateau at lead hour 16 with an MSE of 2.2, hinting towards the model finding a mean value minimizing the forecast error. This can be seen in the MLP prediction flattening out with higher lead times (Figure 8).
Figure 7: Total MSE of the different methods shown for each MADIS station, summed over the lead times. The weather stations along the coast show the highest error and are also where the MPNN + ERA5 model results in the largest improvements over other methods. The station circled in red is the station with the worst MSE for the MPNN + ERA5 model, for which a snippet of time series is shown in Figure 8.
Overall, the model reduces the MSE by 83% compared to the interpolated ERA5, and 47% compared to the MADIS persistence. Spatially, the model improves most along the coast compared to the persistence, better predicting the higher average wind speeds, but also improves on individual stations inland.
@@
Using the multi-modal MPNN to correct the global forecast results in a stark improvement in performance compared to the MLP. The MSE is below 1 $m^{2}/s^{2}$ across all lead times and is reduced by over 50% relative to the MLP. Throughout the different lead times, the MPNN + ERA5 model seems to reach a plateau much faster (around lead hour 8), and at a much lower MSE (0.8 $m^{2}/s^{2}$), demonstrating a clearly better predictive/corrective power. Along the coast is also where the MPNN + ERA5 method shows the strongest improvement in performance: most stations with high errors in the other methods are modeled with a relatively low MSE by the MPNN.
Interpolated ERA5 forecast is of negligible benefit to a simple MLP
In our ablation experiments where we remove ERA5 data from the model, the MLP model performs very similarly. In other words, a spatially-interpolated global forecast does not significantly help the MLP to forecast station-level wind. Across the lead times, all MSE differences are below 0.1 $m^{2}/s^{2}$ between the MLP with and without ERA5.
Message-passing of ERA5 data is crucial to MPNN forecast ability
By contrast, there is a large difference between the MPNN with and without ERA5. The MPNN without ERA5 produces an MSE curve similar to the MLP (maximum difference of 0.3 $m^{2}/s^{2}$ MSE at lead hour 8),
Figure 8: Example time series from the weather station where MPNN + ERA5 performed worst over all lead times. The same station is circled in red on the map in Figure 7. For each experiment, the $u$ and $v$ components are shown separately. Each panel has the MADIS ground truth in blue, and the predictions at increasing lead times displayed with decreasing saturation. Interestingly, even though for this station and time snippet the interpolated ERA5 appears relatively accurate, the MLP experiment does not perform significantly better in correction mode (with ERA5 data) than forecasting mode (no ERA5). In contrast, the MPNN is able to take advantage of the ERA5 data. See Figures A.5 and A.6 for further examples of time series, and Figure A.4 for an example of the station’s environment.
but the introduction of ERA5 significantly improves the model performance. Total MSE of the MPNN without ERA5 is about twice that of the MPNN + ERA5 (13.1 vs 6.0 $m^{2}/s^{2}$). Figure 8 shows, for the different model, a snippet of time series of wind $u$ and $v$ for the station with the worst MSE for the MPNN + ERA5 model. Unlike at many other stations, the interpolated ERA5 data for this location agrees with the local MADIS data well, allowing us to evaluate its integration. The MPNN integrates the ERA5 data the best, consistently correcting it towards the local prediction. The further out the lead time, the more it relies on the ERA5 data, but still performs better than the interpolated ERA5. Meanwhile, the MLP fails to integrate the interpolated ERA5 data, trending towards a flat wind vector with increasing lead times, preferring to minimize error by predicting the average.
@@ 5 Discussion
Discrepancies arise between weather station observations and reanalysis data, such as ERA5, due to the latter’s limitations in modeling terrain and land cover. These local surface characteristics significantly influence local weather patterns, particularly wind, but are not adequately represented in gridded global datasets. The high MSE associated with MADIS persistence underscores the inherent chaotic nature of wind dynamics; current wind conditions do not reliably indicate future conditions. However, this unpredictability has a less pronounced effect on MSE at low wind speeds, typically observed inland and in sheltered coastal regions. Given these complexities, accurate wind forecasting requires a model that accounts for both the discrepancies between global and local conditions and the inherent variability of wind across space and time.
@@
ERA5 and historical station data provide complementary information on wind dynamics. We thus in this work seek a method that can integrate these two datasets successfully for forecasting. ERA5 offers large-scale atmospheric patterns, but it fails to capture low wind speeds in areas that are sheltered from wind due to topography or land surface characteristics. Machine learning methods trained on station data can adapt to local conditions like low wind speeds, but struggle to account for changes between low and high wind speeds (i.e., ramp-ups and ramp-downs) and do not effectively capture the dynamics associated with longer lead times. When predicting local weather using only local data (i.e., in forecasting mode), the MLP and MPNN achieve comparable performance. However, with the introduction of ERA5, the MLP appears to struggle to effectively integrate the interpolated global data with local station observations. In contrast, our proposed approach, MPNN + ERA5, successfully incorporates large-scale atmospheric dynamics from ERA5, thereby improving predictions at longer lead times. This approach also effectively distinguishes among nearby stations with low and high average wind speeds, particularly along the coast where all other models fail. This superior performance suggests that the spatial structure of the GNN, along with the proposed heterogeneous message passing structure, is highly suitable for correcting a globally gridded weather dataset to reflect local conditions. Interestingly, the inductive bias of the station’s spatial correlation within the MPNN does not, on its own, appear to significantly enhance forecasting performance. This may be attributed to insufficient station density, limiting the information that can be inferred from surrounding MADIS nodes.
@@
This study demonstrates the effectiveness of GNNs for integrating diverse data formats in local weather prediction. The proposed GNN model successfully combines the evenly spaced mesh grid of ERA5 data with the irregularly scattered off-grid points of MADIS data.
@@ 6 Conclusion
This work demonstrates the effectiveness of a multi-modal GNN for downscaling gridded weather forecasts and improving the accuracy of off-grid predictions. Our model addresses the inherent bias in gridded reanalysis products like ERA5, which limits the accuracy of traditional ML models for off-grid locations. By incorporating both ERA5 and MADIS weather station data within a heterogeneous graph, our GNN predicts off-grid weather conditions by leveraging both large-scale atmospheric dynamics and local weather patterns. In our evaluation using a surface wind prediction task, the GNN significantly outperformed all baseline models. For instance, it achieved a 92.18% reduction in MSE compared to ERA5 interpolation and a 55.22% improvement over the best-performing multi-Layer perceptron (MLP). An ablation study, where ERA5 input was removed, resulted in an MSE increase from 6.0 to 13.1 $m^{2}/s^{2}$, highlighting the importance of incorporating global atmospheric dynamics for accurate local predictions. This finding motivates the exploration of additional modalities, such as radar measurements and satellite imagery, which could further enhance local forecast accuracy. This research has significant implications for improving weather forecasting, particularly in high-value regions where weather stations can be installed. More accurate off-grid predictions can enhance weather-dependent decision-making in various sectors, including agriculture, wildfire management, transportation, and renewable energy. Future work will focus on expanding the study area and exploring the integration of our GNN model with weather foundation models.
@@ A.2 Model Architecture
Figure A.2: MLP model architecture. (a) MLP model architecture without ERA5 input. Time series of weather variables at a MADIS station are embedded with a linear layer. The embeded features are processed by two linear layers, and finally decoded with another linear layer as prediction. (b) MLP model architecture with ERA5 input. It follows a similar architecture to (a) but with additional inputs of interpolated ERA5 to a MADIS station.
@@ A.4 Stations Environment
Figure A.4: Examples of environments for three selected weather stations. (a) Overview of the three stations. The three selected weather stations are: (b) the station with the worst overall interpolated ERA5 MSE (Figure A.5), (c) station with the best overall interpolated ERA5 MSE (Figure A.6) and (d) station with the worst overall MPNN + ERA5 MSE (Figure 8). (b), (c) and (d) show a zoomed-out view of each area; (b’), (c’) and (d’) are zoomed in. (b) and (d) show good examples of weather stations being surrounded by trees and buildings, where wind is affected by these local surface characteristics.
@@ A.5 Additional Results
Table A.4: Test set MSE of experiments for each lead time. MSE is averaged over stations and time steps for the year 2023.
Model ERA5? Lead Time [hrs] Total MSE
1 2 4 8 16 24 36 48
Interpolated ERA5 9.6 9.6 9.6 9.6 9.6 9.6 9.6 9.6 76.8
Persistence 0.5 0.9 1.5 2.7 4.0 4.6 5.6 5.5 25.2
MLP 0.4 0.7 1.1 1.7 2.2 2.4 2.6 2.6 13.7
MLP + ERA5 0.4 0.7 1.1 1.6 2.2 2.4 2.5 2.6 13.4
MPNN 0.4 0.6 0.9 1.4 2.1 2.4 2.6 2.7 13.1
MPNN + ERA5 0.4 0.5 0.7 0.8 0.9 0.9 0.9 0.9 6.0

### v2

@@ 1 Introduction
  • We evaluate our GNN against a range of data-driven and non-data-driven off-grid weather forecasting methods. Amongst those, our model achieves the best performance. It decreases the average error by 3.57% comparing to the best performing MLP model which reduces the mean wind vector error by 79.34% comparing to interpolated ERA5.
  • We conducted an ablation of ERA5 inputs and observed that a GNN with ERA5 nodes achieves 75% of the error of a GNN without ERA5, indicating that—even in the presence of historical station data—global atmospheric dynamics inform local weather patterns.
@@ 2 Related Work
Weather forecasting has long been a challenging problem in atmospheric sciences, with efforts dating back centuries. Since the advent of numerical weather prediction (NWP) in the mid-20th century, most forecast simulations have been conducted on a regular grid, dividing the atmosphere into evenly spaced discrete points to solve complex partial differential equations. This grid-based approach has remained the foundation of many numerical weather forecasting models such as the Integrated Forecast System (ECMWF, 2022) and High-Resolution Rapid Refresh (Dowell et al., 2022). In recent years, machine learning (ML) has gained traction as a promising tool in weather forecasting (Bauer et al., 2015), offering new techniques to improve accuracy and computational efficiency. These ML weather models can be roughly divided into two categories: end-to-end models and foundation models. FourCastNet (Pathak et al., 2022), GraphCast (Lam et al., 2023), Pangu-Weather (Bi et al., 2023), AIFS (Lang et al., 2024) and NeuralGCM (Kochkov et al., 2024) are end-to-end models trained directly to make weather forecasts. In contrast, AtmoRep (Lessig et al., 2023), ClimaX (Nguyen et al., 2023), Aurora (Bodnar et al., 2024) and Prithvi WxC (Schmude et al., 2024) are foundation models that are first trained with a self-supervision task and then fine-tuned for weather forecasting. However, the training data for ML models largely stem from traditional gridded numerical simulations such as ERA5 (Hersbach et al., 2020). As a result, the ML models themselves still typically maintain the grid-based paradigm even within their more modern forecasting approach. One major disadvantage of gridded weather forecasting is that it is usually limited by its fixed resolution such that it cannot accurately reflect fine-grained local weather patterns (although efforts towards limited area modeling have recently been made (Oskarsson et al., 2023)). Other works focusing on increasing the forecast resolution (Harder et al., 2023; Yang et al., 2023; Prasad et al., 2024) exist, but their methods are mostly tested on synthetic datasets. Work meant to correct ERA5 forecasts exists (Mouatadid et al., 2023), but focus on sub-seasonal forecast at coarse spatial resolution rather than local weather forecasting. In this work, we propose a multi-modal graph neural network which can effectively downscale gridded weather forecasting to match real-world local weather dynamics.
@@ 4 Results
We report the model performance for the 2023 test set as the mean error (ME) of the predicted wind vector against the actual wind vector ($\nicefrac{{1}}{{N}}\sum_{i}^{N}\sqrt{\Delta u_{i}+\Delta v_{i}}$). Figure 6 and Table A.3 summarize the ME for the different lead times and models. The average ME for each experiment (mean across lead times) is also reported in Table A.3.
@@
Interpolated ERA5 data does not describe the local wind conditions accurately, with an ME at all lead times of 2.71 $m/s$. (Note that the ME for interpolated ERA5 is constant across lead times, because the “future” values of ERA5 are assumed to be known. This is the best case for a gridded NWP forecast or an ML weather model trained to predict ERA5; if ERA5 were replaced by the output of a global forecasting model, the error would increase with lead time.) Spatially, when considering the ME averaged over lead times for each station (Figure 7), ERA5 appears to consistently misrepresent local weather patterns. The largest magnitude errors are generally concentrated along the coast, due to the higher average wind speeds there, but the highest relative errors are inland.
Figure 6: Mean error (ME) of the wind vector for different models calculated over all test set time steps and stations, from 1 hour to 48 hours lead times. The interpolated ERA5 error is calculated by linearly interpolating ERA5 to each weather station; it is the error between the global and local dataset. The persistence is calculated by offsetting the ground truth (MADIS) by the lead time. The results for the multi-layer perceptron (MLP) and message passing neural network (MPNN) are reported with and without the inclusion of ERA5 as input. All methods outperform the ERA5 interpolation, all ML methods outperform the persistence, and the MPNN + ERA5 outperforms all other methods.
The persistence baseline shows a low ME for short lead times—lead times 1, 2, and 4 hours all have MEs below 0.75 $m/s$. This is expected due to temporal auto-correlation. ME rapidly increases with longer lead times, reaching a maximum ME of 1.46 $m/s$ at 36 hours. Hour 48 shows a slight decrease in ME, hinting at some daily periodicity in wind patterns. Spatially, most of the high ME is also concentrated along the coast, with some exceptions inland. The ME of inland stations is on average substantially lower for the persistence than the interpolated ERA5 baseline.
@@
The MLP + ERA5 method significantly improves on the non-ML methods, showing an ME below 0.65 $m/s$ for all lead times. The MLP error increases with lead hours as expected, reaching the beginning of a plateau at lead hour 16 with an ME of 0.62, hinting towards the model finding a mean value minimizing the forecast error. This can be seen in the MLP prediction flattening out with higher lead times (Figure 8). Overall, the model reduces the ME by 79% compared to the interpolated ERA5, and 44% compared to the MADIS persistence. Spatially, the model improves most along the coast compared to the persistence, better predicting the higher average wind speeds, but also improves on individual stations inland.
Figure 7: Average ME of the different methods shown for each MADIS station, averaged over the lead times. The weather stations along the coast show the highest error and are also where the MPNN + ERA5 model results in the largest improvements over other methods. The station circled in red is the station with the worst ME for the MPNN + ERA5 model, for which a snippet of time series is shown in Figure 8.
@@
Using the multi-modal MPNN to correct the global forecast results in an improvement in performance compared to the MLP. The ME is below 0.62 $m/s$ across all lead times and is reduced by 3.6% relative to the MLP. Throughout the different lead times, the MPNN + ERA5 model seems to reach a plateau slightly faster (around lead hour 8), and at a lower ME (0.57 $m/s$), demonstrating a better predictive/corrective power. Along the coast is where the MPNN + ERA5 method shows the strongest improvement in performance: most stations with high errors in the other methods are modeled with a relatively low ME by the MPNN.
Correcting ERA5 data appears significantly less complex than forecasting with a weather station’s historical data.
In our ablation experiments where we remove ERA5 data from the model, both the MLP model and the MPNN perform significantly worse. Both MLP and MPNN show an increase in ME of $\sim$33on average, highlighting the significantly harder task of forecasting from the station historical data, compared to correcting the ERA5 data. Figure 8 shows, for the different model, a snippet of time series of wind $u$ and $v$ for the station with the worst ME for the MPNN + ERA5 model. Unlike at many other stations, the interpolated ERA5 data for this location agrees with the local MADIS data well, allowing us to evaluate its integration. The MPNN and MLP model integrate the ERA5 data well, consistently correcting it towards the local prediction. The further out the lead time, the more it relies on the ERA5 data, but still performs better than the interpolated ERA5. Meanwhile, without ERA5, the models fails for longer lead times, trending towards a flat wind vector, preferring to minimize error by predicting the average.
Message-passing improves forecasting ability.
The graph structure improves the ME both with and without ERA5. The introduction of the spatial structure improves the ME by 2.7 and 3.6% (without and with ERA5 respectively). The network density allows neighbors to forward information deeper into the graph.
@@ 5 Discussion
Discrepancies arise between weather station observations and reanalysis data, such as ERA5, due to the latter’s limitations in modeling terrain and land cover. These local surface characteristics significantly influence local weather patterns, particularly wind, but are not adequately represented in gridded global datasets. The high ME associated with MADIS persistence underscores the inherent chaotic nature of wind dynamics; current wind conditions do not reliably indicate future conditions. However, this unpredictability has a less pronounced effect on ME at low wind speeds, typically observed inland and in sheltered coastal regions. Given these complexities, accurate wind forecasting requires a model that accounts for both the discrepancies between global and local conditions and the inherent variability of wind across space and time. ERA5 and historical station data provide complementary information on wind dynamics. We thus in this work seek a method that can integrate these two datasets successfully for forecasting. ERA5 offers large-scale atmospheric patterns, but it fails to capture low wind speeds in areas that are sheltered from wind due to topography or land surface characteristics.
@@
Machine learning methods trained on station data can adapt to local conditions like low wind speeds, but struggle to account for changes between low and high wind speeds (i.e., ramp-ups and ramp-downs) and do not effectively capture the dynamics associated with longer lead times. When predicting local weather using only local data (i.e., in forecasting mode), the MLP and MPNN achieve respectable performances. However, with the introduction of ERA5, the MLP and MPNN, both successfully incorporates large-scale atmospheric dynamics from ERA5, thereby improving predictions at longer lead times. The slightly superior performance of the MPNN suggests that the spatial structure of the GNN, along with the proposed heterogeneous message passing structure, can be suitable for correcting a globally gridded weather dataset to reflect local conditions. Even without ERA5, the inductive bias of the station’s spatial correlation within the MPNN does, on its own, enhance forecasting performance.
@@
This study demonstrates the effectiveness of ML methods for integrating diverse data formats in local weather prediction. The proposed GNN model successfully combines the evenly spaced mesh grid of ERA5 data with the irregularly scattered off-grid points of MADIS data.
Figure 8: Example time series from the weather station where MPNN + ERA5 performed worst over all lead times. The same station is circled in red on the map in Figure 7. For each experiment, the $u$ and $v$ components are shown separately. Each panel has the MADIS ground truth in blue, and the predictions at increasing lead times displayed with decreasing saturation. Interestingly, for this station and time snippet the interpolated ERA5 appears relatively accurate, and the MLP and MPNN are able to take advantage of the ERA5 data. See Figures A.4 and A.5 for further examples of time series, and Figure A.3 for an example of the station’s environment.
@@ 6 Conclusion
This work demonstrates the use of a multi-modal GNN for downscaling gridded weather forecasts and improving the accuracy of off-grid predictions. Our model addresses the inherent bias in gridded reanalysis products like ERA5. By incorporating both ERA5 and MADIS weather station data within a heterogeneous graph, our GNN predicts off-grid weather conditions by leveraging both large-scale atmospheric dynamics and local weather patterns. In our evaluation using a surface wind prediction task, the GNN outperformed all baseline models. For instance, it achieved a 80% reduction in ME compared to ERA5 interpolation and a 3.6% improvement over the best-performing multi-Layer perceptron (MLP). An ablation study, where ERA5 input was removed, resulted in an ME increase from 0.54 to 0.72 $m/s$, highlighting the importance of incorporating global atmospheric dynamics for accurate local predictions. This finding motivates the exploration of additional modalities, such as radar measurements and satellite imagery, which could further enhance local forecast accuracy. This research has significant implications for improving weather forecasting, particularly in high-value regions where weather stations can be installed. More accurate off-grid predictions can enhance weather-dependent decision-making in various sectors, including agriculture, wildfire management, transportation, and renewable energy. Future work will focus on expanding the study area and exploring the integration of our GNN model with weather foundation models.
@@ A.4 Stations Environment
Figure A.3: Examples of environments for three selected weather stations. (a) Overview of the three stations. The three selected weather stations are: (b) the station with the worst overall interpolated ERA5 ME (Figure A.4), (c) station with the best overall interpolated ERA5 ME (Figure A.5) and (d) station with the worst overall MPNN + ERA5 ME (Figure 8). (b), (c) and (d) show a zoomed-out view of each area; (b’), (c’) and (d’) are zoomed in. (b) and (d) show good examples of weather stations being surrounded by trees and buildings, where wind is affected by these local surface characteristics.
@@ A.5 Additional Results
Table A.3: Test set Wind Vector Error of experiments for each lead time. Wind Vector Error is averaged over stations and time steps for the year 2023.
Model ERA5? Lead Time [hrs] Average
1 2 4 8 16 24 36 48 ME
Interpolated ERA5 2.71 2.71 2.71 2.71 2.71 2.71 2.71 2.71 2.71
Persistence 0.41 0.55 0.74 1.00 1.22 1.25 1.46 1.38 1.00
MLP 0.38 0.5 0.64 0.78 0.87 0.9 0.92 0.92 0.74
MLP + ERA5 0.38 0.47 0.54 0.59 0.62 0.63 0.64 0.64 0.56
MPNN 0.37 0.47 0.59 0.71 0.83 0.9 0.93 0.93 0.72
MPNN + ERA5 0.37 0.45 0.53 0.57 0.59 0.59 0.6 0.62 0.54